# Optimizing Calibration by Gaining Aware of Prediction Correctness

### v2

@@ 3.3 Gaining correctness awareness
Our solution to gaining correctness awareness is motivated by (Deng et al., 2022). They find consistency of model predictions of transformed (e.g., rotated, gray-scaled, etc.) images is highly correlated with model accuracy. While the insight from (Deng et al., 2022) is on the dataset level, our assumption is on the individual sample level: model behavior on transformed samples is useful to inform the correctness of model predictions.
The pipeline of our calibrator is presented in Fig. 2, we aim to calibrate a classification ‘model’ $f$. To do so, we compute the logit vector $\mathbf{z}$, softmax vectors of an original image $\mathbf{X}$ and its transformed versions $\mathbf{v}$ and $\mathbf{v}_{i}$ ($i\in\mathcal{I}_{M}$, assuming $M$ types of transforms), respectively. We determine the indices $\mathbf{q}\in\mathbb{R}^{k}$ of the $k$ largest softmax scores of $\mathbf{v}$, and use these indices $\mathbf{q}$ to locate and select the corresponding values from $\mathbf{v}_{i}$, forming new vector with $k$ dimensions. These $k$-dim vectors $\mathbf{v}_{i}[\mathbf{q}]$ of the transformed images are concatenated as $\oplus_{i\in\mathcal{I}_{M}}\mathbf{v}_{i}[\mathbf{q}]$, and used as the calibrator input.
Calibrator, $g$ parameterized by $\theta$, which is trained by the proposed CA loss. Here, the calibrator consists of two fully connected layers with a ReLU activation function in between. Each hidden layer comprises 5 nodes. The calibrator is optimized on the calibration (a.k.a., validation) set, and the calibration output is temperature $\tau$ to be used to scale the model logit vector $\mathbf{z}$ of original image. Below we show these steps in equations.
$\mathbf{q}=\underset{\mathbf{q}}{\text{argmax}}\mathbf{v}\in\mathbb{R}^{k}$ (16)
$\tau=g_{\theta}(\oplus_{i\in\mathcal{I}_{M}}\mathbf{v}_{i}[\mathbf{q}])$ (17)
@@
$\arg\min_{\theta}\mathbb{E}_{f}^{\text{emp}}(\hat{c},y_{\mathbf{x}},\hat{y}_{\mathbf{x}}),$ (19)
where $y_{\mathbf{x}}$ and $\hat{y}_{\mathbf{x}}$ denote respectively the ground truth and predicted labels. Algorithm 1 summarises this calibration process.
Algorithm 1 Calibrator with our proposed Correctness-Aware Loss
  Return: Calibrator model weights $\theta$.
In practice, transformations can be grayscale, rotation, color jitter, adding Gaussian noise, random erasing, etc. They are applied to the original image during training and inference. We do not assume access to test batches, which are used in some previous works (Guo et al., 2017; Wang et al., 2023). Nevertheless, if we assume such access, we can retrieve from test batches images that are similar to the original one and use softmax vectors of the retrieved images as calibrator inputs. As to be shown in Sec. 4.4, we find that grayscale, rotation, and colorjitter are effective ones, and that using four transformations give a good trade-off between calibration performance and computational cost.
During inference, using the $k$-dim vectors from the transformed images, we obtain an adjusted temperature from the calibrator. This temperature is used to scale the logits of the original image, the softmax vector of which is then updated.
Figure 3: Comparison of different loss functions w.r.t. temperature and softmax probability of the ground truth (GT) class. In a four-way classification task, we examine a wrongly predicted sample with logit vector $[a,2.0,0.1,0.05]$, where $a<2$ is the value on the ground truth class. We use $c_{\text{gt}}$ to donate the softmax score of the GT class. Top: The loss surface plots for varying temperatures and $c_{\text{gt}}$, with red and blue arrows representing positive and negative temperature gradients, respectively. Bottom: Shows 2D loss curves for varying $c_{\text{gt}}$. The lines in the bottom charts correspond to the lines of the same color in the top charts. Compared with traditional Maximum Likelihood Estimation (MLE) based functions (e.g., Cross-Entropy, Mean Squared Error), our Correctness-Aware loss minimization does not favor temperatures below 1 for incorrect predictions, while sometimes MLE does.
@@ 3.4 Comparison between CA loss and MLE
Maximum Likelihood Estimate ( MLE) is widely used for calibration training (Kumar et al., 2019), under concrete formats such as the Cross-Entropy (CE) or Mean Square Error (MSE) losses. This section goes through their connections and differences with the CA loss.
Empirically, we find that such narrowly wrong predictions take up 2%-8% of the calibration set (ImageNet validation) [^4] We first compute the ratio of the probability on the ground-truth class to that on the wrongly predicted class. We define a sample is narrowly wrong prediction if this ratio is higher than 0.5.. This would negatively impact training efficacy of MLE. Moreover, during inference, if a test set has many such narrowly wrong predictions, MLE will also be negatively impacted because of its unsuitable in dealing with such samples during training. This would explain why our system is superior to and on par with state of the art on OOD and IND test sets, respectively (refer to Sec. 4). But in both IND and OOD scenarios, our calibrated models are much better than uncalibrated ones.
Table 1: Calibrator comparison under the ImageNet setup. Each reported number is averaged over 10 classifiers, described in Section 4.1. We use five test sets: ImageNet-Val, ImageNet-A, ImageNet-R, ImageNet-S, and ObjectNet, and four metrics: ECE (bin=25), BS, KS and AUC (AUROC). Best results in each column are in bold. When comparing CA and CE, better results are highlighted in blue.
@@
3.5 Discussion
Table 2: Calibrator comparison under the CIFAR-10 setup. Each number is averaged over 10 classifiers (see Section 4.1). We use one IND test set (CIFAR10.1) and three OOD test sets (CIFAR-10.1 , CINIC, and CIFAR-10-C). Other notations are the same as Table 1.
@@ 3.5 Discussion
Correctness prediction performance as a potential calibration metric. Given a confidence value, better separability between correct and wrong predictions leads to safer decision making for users, because less mistakes are made. This paper uses area under the ROC curve (AUROC) to measure the performance of predicting classification correctness. As shown in Section B.1 of Appendix, a small ECE does not always mean user makes less mistakes during decision making. Considering the strong tie between this 2-way classification problem and model calibration (Eq. (1)), we think AUROC can be an additional evaluation metric for model calibration. In Section B.1 of Appendix, more comparisons between AUROC and ECE are presented.
@@ 4.1 Models and datasets
iWildCam setup. 1. Model. We use 10 models trained on the iWildCam(Beery et al., 2020) training set. They are downloaded from the official dataset website. 2. Calibration set. We train the calibrator on the iWildCam validation set. 3. Test set. We use the iWildCam test set containing animal pictures captured in the wild. Further details of the three setups are provided in Sec. A of Appendix.
@@ 4.2 Calibration methods and evaluation metrics
Methods. We compare our method with six popular calibration methods. They include scaling-based methods such as temperature scaling (TS) (Guo et al., 2017), ensemble temperature scaling (ETS) (Zhang et al., 2020), adaptive temperature scaling (Adaptive TS) (Joy et al., 2023), and parameterized temperature scaling (PTS) (Tomani et al., 2022). We also compare with a binning method multi-isotonic regression (MIR) (Zhang et al., 2020), True Class Probability (TCP) (Corbiere et al., 2021), a spline-based re-calibration method (Spline) (Gupta et al., 2020), and Proximity-Informed Calibration (ProCal) (Xiong et al., 2024).
Metrics. Apart from the expected calibration error (ECE) (Guo et al., 2017), we report the Brier score (BS), adaptive calibration error (ACE) (Nixon et al., 2019), and Kolmogorov-Smirnov (KS) error (Gupta et al., 2020). In addition, we use area under the ROC curve (AUROC) to evaluate how well the calibrators separate correct predictions from wrong predictions, which, as mentioned in Sec. 3.5, might also be a good metric for calibration. All the numbers in Table 12, and 3 are averaged over results of calibrating 10 different models introduced in Sec. 4.1.
@@ 4.3 Main observations
Comparison of calibration performance with the state of the art. We summarize calibration results under the ImageNet, CIFAR-10, and iWildCam setups in Table 1, Table 2, and Table 3, respectively. We have two observations. First, on OOD test sets, our method is very competitive across various metrics. For example, when compared with the second-best method on ObjectNet, the ECE, BS, and KS metrics of our method are 10.56%, 3.69%, and 8.97% lower, respectively. Second, on near IND or IND test sets such as ImageNet-Val and CIFAR-10.1 , our method is less advantageous but is still competitive. The reason for our method being more effective on OOD test sets is that there are more narrowly wrong predictions, mentioned in the last paragraph in Section 3.4. Besides, as explained in Section 3.3, the use of transformed images might not be an optimal way to inform classification correctness.
Comparing CA with MLE. In Table 1, Table 2, and Table 3, we compare ‘CA only’ with ‘CE only’, and ‘CA+trans’ with ‘CE+trans’. First, ‘CA only’ consistently outperforms ‘MSE only’ in 18 out of 20 scenarios under the ImageNet setup, 13 out of 16 scenarios under the CIFAR-10 setup, and 4 out of 4 scenarios under the iWildCam setup. Second, in most cases (e.g., 19 out of 20 scenarios under ImageNet setup), ‘CA+trans’ is better than ‘CE+trans’. In addition, in Table 1, we observe that the combination of CE and CA does not yield better results compared to using CA alone in the OOD test set. The superiority of the CA loss is more evident on OOD datasets as discussed in Sec. 3.5.
Potential of CA in allowing confidence to better separate correct and wrong predictions. In Tables 1, 2, and 3, we compare separability and have two observations. First, existing methods typically do not have improvement in AUROC. This is not surprising, because their working mechanisms are not relevant to the separation of correct and wrong predictions. Second, our method improves AUROC under the ImageNet and iWildCam setups and in on par with existing methods on CIFAR-10. In fact, we find predictions of transformed images offer much less diversity under CIFAR-10 classifiers, losing their efficacy in telling prediction correctness. This could be addressed with a better method than transformed images, and we leave it for future work. These results, especially those under the challenging ImageNet and iWildCam setups, suggest our method has the potential to better distinguish between correct and wrong predictions by confidence scores, which could lead to improved decision making. A closer look at the ROC curves is provided in Fig. 5 and Sec. B.1 of Appendix.
Table 3: Calibrator comparison under iWildCam setup. Each number is averaged over 10 classifiers (see Section 4.1). We use the iWildCam test set. Other notations are the same as Table 1.
@@
Figure 5: (Left:) Comparing various combinations of image transformations, including rotation (R), grayscale (S), colorjitter (C), random erasing (E) and Gaussian noise (G). Different colors means different numbers of transformations. Dashed lines denote performance of no calibration and retrieval-based augmentation that accesses test batches. (Right:) Visualization of ROC curves of various calibrators. Existing methods typically do not improve AUC, while our method effectively does. All results in this figure are reported for ObjectNet using the model ‘beit_base_patch16_384’, as introduced in Appendix A.
Effectiveness of using transformed images as calibrator input. We compare ‘CA+trans’ and ‘CA only’ in Table 1, Table 2, and Table 3. It is very clearly that ‘CA+trans’ gives consistently better calibration performance than ‘CA only’. It indicates the necessity of using transformed images.
4.4 Further analysis
@@
On the other hand, from Fig. 4 (right), when training set contains lots of narrowly wrong predictions, CE is very poor and even worse then no calibration. When more absolutely wrong samples are included, CE becomes gradually better and even close to our method. These results empirically verifies our analysis in Section 3.4.
@@ 5 Conclusion
Appendix A Experimental Setup

### v3

@@ 3.3 Gaining correctness awareness
Our approach, inspired by (Deng et al., 2022), leverages the discovery that consistency in model predictions for transformed images correlates strongly with accuracy. While their insight focuses on dataset-level consistency, our assumption extends to individual samples: the model’s behavior on transformed samples informs prediction correctness.
The pipeline of our calibrator is presented in Fig. 2, we aim to calibrate a classification model $f$. To do so, we compute the logit vector $\mathbf{z}$, softmax vectors of an original image $\mathbf{X}$ and its transformed versions $\mathbf{v}$ and $\mathbf{v}_{i}$ ($i\in\mathcal{I}_{M}$, assuming $M$ types of transforms), respectively. We determine the indices $\mathbf{q}\in\mathbb{R}^{k}$ of the $k$ largest softmax scores of $\mathbf{v}$, and use these indices $\mathbf{q}$ to locate and select the corresponding values from $\mathbf{v}_{i}$, forming new vector with $k$ dimensions. These $k$-dim vectors $\mathbf{v}_{i}[\mathbf{q}]$ of the transformed images are concatenated as $\oplus_{i\in\mathcal{I}_{M}}\mathbf{v}_{i}[\mathbf{q}]$, and used as the calibrator input.
Calibrator, $g$ parameterized by $\theta$, which is trained by the proposed CA loss. Here, the calibrator consists of two fully connected layers with a ReLU activation function in between. Each hidden layer comprises 5 nodes. The calibrator is optimized on the calibration (a.k.a., validation) set, and the calibration output is temperature $\tau$ to be used to scale the model logit vector $\mathbf{z}$ of original image. Below we show these steps in equations:
$\mathbf{q}=\underset{\mathbf{q}}{\text{argmax}}\mathbf{v}\in\mathbb{R}^{k},$ (16)
$\tau=g_{\theta}(\oplus_{i\in\mathcal{I}_{M}}\mathbf{v}_{i}[\mathbf{q}]),$ (17)
@@
$\arg\min_{\theta}\mathcal{L}_{\text{CA}}(\hat{c},y_{\mathbf{x}},\hat{y}_{\mathbf{x}}),$ (19)
where $y_{\mathbf{x}}$ and $\hat{y}_{\mathbf{x}}$ denote respectively the ground truth and predicted labels.
In practice, transformations can be grayscale, rotation, color jitter, adding Gaussian noise, random erasing, etc. They are applied to the original image during training and inference. We do not assume access to test batches, which are used in some previous works (Guo et al., 2017; Wang et al., 2023). Nevertheless, if we assume such access, we can retrieve from test batches images that are similar to the original one and use softmax vectors of the retrieved images as calibrator inputs. As to be shown in Sec. 5, we find that grayscale, rotation, and colorjitter are an effective combination, and that using three transformations give a good trade-off between calibration performance and computational cost.
During inference, using the $k$-dim vectors from the transformed images, we obtain an adjusted temperature from the calibrator. Alg. 1 in Appendix A summarises this calibration process.
@@ 3.4 Comparison between CA loss and MLE
When a sample is correctly predicted, both CA loss and MLE (e.g., cross-entropy, mean squared error) behave similarly, pushing the ground-truth probability toward 1 and thereby improving calibration. However, for wrong predictions, MLE can inadvertently raise confidence for narrowly wrong samples. As shown in Fig. 3, for a narrowly wrong prediction (yellow curve), MSE and Cross-Entropy minimize their loss at a temperature greater than 1—contradicting the goal of lowering confidence for incorrect predictions. In contrast, CA loss steadily decreases confidence in such cases, resulting in better overall calibration. Please see Appendix B for more detailed discussion.
Figure 3: Comparison of different loss functions w.r.t. temperature and softmax probability of the ground truth (GT) class. In a four-way classification task, we examine a wrongly predicted sample with logit vector $[a,2.0,0.1,0.05]$, where $a\!<\!2$ is the value on the ground truth class. We use $c_{\text{gt}}$ to donate the softmax score of the GT class. Top: The loss surface plots for varying temperatures and $c_{\text{gt}}$, with red and blue arrows representing positive and negative temperature gradients, respectively. Bottom: Shows 2D loss curves for varying $c_{\text{gt}}$. The lines in the bottom charts correspond to the lines of the same color in the top charts. Compared with Maximum Likelihood Estimation (MLE) based functions (e.g., Cross-Entropy, Mean Squared Error), our Correctness-Aware loss minimization does not favor temperatures below 1 for incorrect predictions, while sometimes MLE does.
Table 1: Calibrator comparison under the ImageNet setup. Each reported number is averaged over 10 classifiers, described in Sec. 4.1. We use five test sets: ImageNet-Val, ImageNet-A, ImageNet-R, ImageNet-S, and ObjectNet, and four metrics: ECE (bin=25), BS, KS and AUC (AUROC). Best results in each column are in bold. When comparing CA and CE, better results are in blue color.
@@
Table 2: Calibrator comparison under the CIFAR-10 setup. Each number is averaged over 10 classifiers (see Sec. 4.1). We use one IND test set (CIFAR10.1) and three OOD test sets (CIFAR-10.1 , CINIC, and CIFAR-10-C). Other notations are the same as Table 1.
@@ 4.1 Models and datasets
iWildCam setup. 1. Model. We use 10 models trained on the iWildCam(Beery et al., 2020) training set. They are downloaded from the official dataset website. 2. Calibration set. We train the calibrator on the iWildCam validation set. 3. Test set. We use the iWildCam test set containing animal pictures captured in the wild. Further details of the three setups are provided in Appendix E.
@@ 4.2 Calibration methods and evaluation metrics
Methods. We compare our method with 8 popular calibration methods. They include scaling-based methods such as temperature scaling (TS) (Guo et al., 2017), ensemble temperature scaling (ETS) (Zhang et al., 2020), adaptive temperature scaling (Adaptive TS) (Joy et al., 2023), and parameterized temperature scaling (PTS) (Tomani et al., 2022). We also compare with binning method multi-isotonic regression (MIR) (Zhang et al., 2020), True Class Probability (TCP) (Corbiere et al., 2021), spline-based re-calibration method (Spline) (Gupta et al., 2020), and Proximity-Informed Calibration (ProCal) (Xiong et al., 2024).
Metrics. Apart from the expected calibration error (ECE) (Guo et al., 2017), we report the Brier score (BS), adaptive calibration error (ACE) (Nixon et al., 2019), and Kolmogorov-Smirnov (KS) error (Gupta et al., 2020). In addition, we use area under the ROC curve (AUROC) to evaluate how well the calibrators separate correct predictions from wrong predictions, which might also be a good metric for calibration (Appendix D). All the numbers in Table 12, and 3 are averaged over results of calibrating 10 different classification models (Appendix E).
@@ 4.3 Main observations
Table 3: Calibrator comparison under iWildCam setup. Each number is averaged over 10 classifiers. We use the iWildCam test set. Other notations are the same as Table 1.
@@
Comparison of calibration performance with the state of the art. We summarize calibration results under the ImageNet, CIFAR-10, and iWildCam setups in Table 1, 2, and 3, respectively. We have two observations. First, on OOD test sets, our method is very competitive across various metrics. For example, when compared with the second-best method on ObjectNet, the ECE, BS, and KS metrics of our method are 10.56%, 3.69%, and 8.97% lower, respectively. Second, on near IND or IND test sets such as ImageNet-Val and CIFAR-10.1 , our method is less advantageous but is still competitive. The reason for our method being more effective on OOD test sets is that there are more narrowly wrong predictions, mentioned in the last paragraph in Sec. B. Besides, as explained in Sec. 3.3, the use of transformed images might not be an optimal way to inform classification correctness.
Comparing CA with MLE. In Table 1, 2, and 3, we compare ‘CA only’ with ‘CE only’, and ‘CA+trans’444‘CA only’ uses the CA loss and the top-$k$ logits of the original image as inputs to train the calibrator. ‘CA+trans’ uses the CA loss and logits from transformed images instead (see, Fig. 2). with ‘CE+trans’. First, ‘CA only’ consistently outperforms ‘MSE only’ in 18 out of 20 scenarios under the ImageNet setup, 13 out of 16 scenarios under the CIFAR-10 setup, and 4 out of 4 scenarios under the iWildCam setup. Second, in most cases (e.g., 19 out of 20 scenarios under ImageNet setup), ‘CA+trans’ is better than ‘CE+trans’. In addition, in Table 1, we observe that the combination of CE and CA does not yield better results compared to using CA alone in the OOD test set. The superiority of the CA loss is more evident on OOD datasets as discussed in Appendix D.
Potential of CA in allowing confidence to better separate correct and wrong predictions. In Tables 1, 2, and 3, we compare separability and have two observations. First, existing methods typically do not have improvement in AUROC. This is not surprising, because their working mechanisms are not relevant to the separation of correct and wrong predictions. Second, our method improves AUROC under the ImageNet and iWildCam setups and in on par with existing methods on CIFAR-10. In fact, we find predictions of transformed images offer much less diversity under CIFAR-10 classifiers, losing their efficacy in telling prediction correctness. This could be addressed with a better method than transformed images, and we leave it for future work. These results, especially those under the challenging ImageNet and iWildCam setups, suggest our method has the potential to better distinguish between correct and wrong predictions by confidence scores, which could lead to improved decision making. A closer look at the ROC curves is provided in Fig. 4 and Appendix F.
Figure 4: (Left:) Comparing various combinations of image transformations, including rotation (R), grayscale (S), colorjitter (C), random erasing (E) and Gaussian noise (G). Different colors means different numbers of transformations. Dashed lines denote performance of no calibration and retrieval-based augmentation that accesses test batches. (Right:) Visualization of ROC curves of various calibrators. Existing methods typically do not improve AUC, while our method effectively does. All results in this figure are reported for ObjectNet using the model ‘beit_base_patch16_384’, as introduced in Appendix E.
@@
Effectiveness of using transformed images as calibrator input. We compare ‘CA+trans’ and ‘CA only’ in Table 1, 2, and 3. It is very clearly that ‘CA+trans’ gives consistently better calibration performance than ‘CA only’. It indicates the necessity of using transformed images.
5 Further analysis
@@
On the other hand, from Fig. 5 (right), when training set contains lots of narrowly wrong predictions, CE is very poor and even worse then no calibration. When more absolutely wrong samples are included, CE becomes gradually better and even close to our method. These results empirically verifies our discussion in Sec. B.
@@ 6 Conclusion
Appendix A Our Algorithm
Alg. 1 shows the calibrator with our proposed Correctness-Aware loss.
Algorithm 1 Calibrator with our proposed Correctness-Aware Loss
  Return: Calibrator model weights $\theta$.
Appendix B Comparison between CA loss and MLE
Maximum Likelihood Estimate ( MLE) is widely used for calibration training (Kumar et al., 2019), under concrete formats such as the Cross-Entropy (CE) or Mean Square Error (MSE) losses. This section goes through their connections and differences with the CA loss. Appendix D shows more discussions.
Empirically, we find that such narrowly wrong predictions take up 2%-8% of the calibration set (ImageNet validation). [^5] We first compute the ratio of the probability on the ground-truth class to that on the wrongly predicted class. We define a sample is narrowly wrong prediction if this ratio is higher than 0.5. This would negatively impact training efficacy of MLE. Moreover, during inference, if a test set has many such narrowly wrong predictions, MLE will also be negatively impacted because of its unsuitable in dealing with such samples during training. This would explain why our system is superior to and on par with state of the art on OOD and IND test sets, respectively (refer to Sec. 4). In both IND and OOD scenarios, our calibrated models are much better.
Appendix C Compared to Pre-hoc Calibration Methods
Although our method focuses on post-hoc model calibration, we compare our method with pre-hoc calibration methods here. We keep the calibrator networks and inputs the same, replacing only the CA loss with the losses from MDCA (Hebbalaguppe et al., 2022) and DCA (Liang et al., 2020). The results are shown in Table 4. We observe that our CA loss outperforms DCA and MDCA.
We also explore potential complementary strengths between train-time calibration methods and our approach. To this end, we combined our CA loss with the loss in MDCA or DCA to train an image transformation-based calibrator. However, we cannot observe complementary strengths of MDCA loss or DCA loss to CA loss in Fig 4. This suggests that some training-time calibration methods may not directly benefit the post-hoc calibration system. They may also face the "narrowly wrong prediction issue".
Table 4: Calibrator comparison under the ImageNet setup. Each reported number is averaged over 10 classifiers. We use four test sets: ImageNet-A, ImageNet-R, ImageNet-S, and ObjectNet, and four metrics: ECE (bin=25), BS, KS, and AUC (AUROC). Best results in each column are in bold.
Method ImageNet-A ImageNet-R ImageNet-S ObjectNet
[…]
Uncal 39.44 32.90 43.47 61.87 13.97 16.89 19.90 88.06 20.92 21.67 29.01 82.57 31.21 25.20 36.48 78.05
DCA 32.10 25.32 36.33 63.33 10.00 14.03 17.99 89.05 17.47 18.00 25.07 84.04 25.25 20.75 31.13 78.54
MDCA 27.80 21.71 32.24 63.32 7.37 13.85 16.28 88.04 13.09 16.97 22.51 82.70 21.11 18.49 28.02 78.05
CA (ours) 20.65 16.79 22.50 63.74 4.91 12.21 10.12 90.22 4.00 13.83 13.12 84.87 10.33 14.59 18.72 79.25
DCA+CA 30.31 24.00 34.56 63.45 8.74 13.50 16.90 89.25 15.29 17.01 23.34 84.32 23.32 19.66 29.51 78.65
MDCA+CA 26.36 20.61 30.63 63.52 6.59 13.41 15.31 88.38 11.18 16.20 21.04 83.10 19.38 17.61 26.57 78.22
Appendix D Further Discussion
Correctness prediction performance as a potential calibration metric. Given a confidence value, better separability between correct and wrong predictions leads to safer decision making for users, because less mistakes are made. This paper uses area under the ROC curve (AUROC) to measure the performance of predicting classification correctness. As shown in Sec. F.2, using a calibration method does not always mean user makes less mistakes during decision making. Considering the strong tie between this 2-way classification problem and model calibration (Eq. (1)), we think AUROC can be an additional evaluation metric for model calibration.
The significance of the bounds of CA loss. The significance of the bounds for the CA loss lies in understanding its behavior during training:
  • Lower bound: Demonstrates whether the loss function can converge during the optimization process, ensuring stability and effectiveness.
  • Upper bound: Indicates the maximum possible value of the loss, useful for diagnosing training stability and ensuring expected behavior.
Appendix E Access Of Benchmarks And Models